OpenStudy (anonymous):

Michelle Payne deposited $8,000 in a savings account paying 6.25% simple interest. How long will it take for her investment to amount to $10,000?

OpenStudy (tkhunny):

I = prt -- Every time. 10000 - 8000 = 2000 = I Go!

OpenStudy (cj49):

Simple interest formula: A=P(1+rt) 10000=8000(1+0.0625t) P=present amount r=annual interest t=number of years

OpenStudy (anonymous):

i have done it like the both of you said and it is still saying I am wrong 8000*(1+.0625) 8000*1.0625=8500... dont get it at all!!

OpenStudy (tkhunny):

You missed cj49's "t". You can't just throw it out. You must find it.

OpenStudy (anonymous):

so i take 8500/10000 right?

OpenStudy (tkhunny):

Don't try to work in pieces. Write the WHOLE equation. 10000=8000(1+0.0625t) Now, divide by 8000 10000/8000 = (1+0.0625t) 1.25 = (1+0.0625t)

OpenStudy (anonymous):

i have done it and i am not getting it ugh business math is for smart ppl not me.

OpenStudy (tkhunny):

You have done what? SHOW your work. We cannot help you if we cannot see what you are doing! 1.25 = (1+0.0625t) Subtract 1 1 = 0.0625t Divide by 0.0625 1/0.0625 = t t = 16 What were you doing that resulted in something else?